Fermi National Laboratory


Accelerator Update for April 1, 2005 - April 4, 2005

Friday April 1
The day shift began with Operations monitoring store 4072, with the Antiproton Source (Pbar) stacking, with MiniBooNE off due to a magnet problem, and with the Meson experiment MCenter taking beam.

At 8:58 AM, Operations halted beam in the Main Injector (MI) to investigate the MiniBooNE magnet HV860. This also halted beam to Pbar and to Switchyard (SY). Linac and Booster took advantage of this downtime and accessed their tunnels for maintenance and repair.

MI experts reported at 11:06 AM, that only the power flag would have to replaced for the E:HV860-2 magnet repair.

Linac came out of access at 11:07 AM.

Booster came out of access and experts began turning on the machine at 12:18 PM.

Experts reported at 1:28 PM that the HV860-2 magnet work was finished.

RF experts accessed Booster at 3:18 PM to bring BRF19 back online. They were out of access by 4:27 PM.

At 4:32 PM, Operations reported a problem with a MI power supply loop controller (MI-60S). Operators called in an expert who had it working by 4:42 PM.

Operations had beam to the Booster beam dump by 4:46 PM.

Pbar resumed stacking at 5:21 PM.

Operations reported at 6:06 PM, that they had established beam to MCenter and MTest.

Operations established beam to NuMI at 6:28 PM so experts could scan their target. They had finished their scans by 9:05 PM.

At 9:21 PM, Operations reported having problems with Pbar stacking and emittances.

MTest experimenters reported at 9:37 PM that they were done with the beam for the night.

Operations reported at 11:53 PM that they resolved the Pbar emittance and stacking problems.

Saturday April 2
The midnight shift began with Operations monitoring store 4072, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 4072 at 12:38 AM.

Operations put the Tevatron (TeV) into shot setup at 12:57 AM.

Operations established store 4073 at 2:59 AM with an initial luminosity of 82.02E30. Pbar resumed stacking.

MCenter resumed taking beam at 3:31 AM.

At 7:05 AM, Operations reported that a TeV RF station (TRF7) tripped off.

At 7:56 AM, a Booster RF station (BRF7) tripped off for the first of five times in thirty minutes. It reset okay each time.

Operations put the Recycler into shot setup to stash antiprotons at 10:19 AM.

At 10:54 AM, BRF16 tripped off for the first of two times in ten minutes.

At 11:03 AM, Operations reported that after TRF7 had tripped off, store 4073's luminosity dropped off drastically. Operations contacted an expert who advised them as to how to stop the decline.

At 11:25 AM, operators responded to a LCW flow problem on the Booster East Anode Power Supply. They couldn't find a problem and by 11:45 AM had restored beam to Pbar and to MCenter.

At 12:11 PM, Operations reported that the Recycler transfers suffered poor efficiency due to emittance problems.

At 12:17 PM, BRF3 & 7 tripped off for the first of two times in thirty minutes. They reset okay.

At 12:35 PM, BRF16 tripped off. It reset okay.

At 12:56 PM, BRF7 tripped off for the first of two times in two minutes. It rest okay, but stations 6, 7, & 9 all started sparking after the last reset.

At 1:01 PM, BRF3, 5, & 7 all tripped off and were reset okay.

At 1:06 PM, Operations reported that they were lowering the Booster intensity to investigate the Booster sparking.

At 1:35 PM, BRF7 tripped off. It reset okay.

At 4:33 PM, MIRF14 tripped off for the first of five times in three and a half hours. It reset okay each time.

At 6:54 PM, BRF7 & 8 tripped off due to high vacuum pressure. They both reset okay.

Operations terminated store 4073 at 8:04 PM. An RF expert began repairs on TRF7. He completed the repair at 8:51 PM.

Operations put the TeV into shot setup at 8:51 PM.

Operations established a combination store 4074 at 11:04 PM with an initial luminosity of 89.21E30.

Pbar and MCenter resumed taking beam at 11:22 PM.

Sunday April 3
The midnight shift began with Operations monitoring store 4074, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the Recycler into shot setup at 7:23 AM.

Recycler stashing ended and Pbar resumed stacking at 8:29 AM and MCenter resumed taking beam nine minutes later.

At 11:29 AM, BRF16 tripped off. It reset okay.

NuMI experts began a target scan at 11:38 AM.

At 12:43 PM, MIRF14 tripped off for the first of two times in two minutes. It reset okay both times.

At 2:33 PM, a Pbar Debuncher power supply (D:H914) tripped off. Operators responded, but couldn't reset the supply. They called in an expert who came in, replaced a fuse, and had the supply back up thirty minutes later.

Pbar resumed stacking at 3:01 PM and MCenter resumed taking beam at 3:09 PM.

Operations put the Recycler into shot setup at 3:32 PM.

The Recycler stash was completed by 4:26 PM. Pbar resumed stacking and MCenter resumed taking beam.

NuMI experts began more target scans at 4:54 PM.

At 5:47 PM, MIRF14 tripped off for the first of two times in thirty minutes.

At 7:46 PM, BRF16 tripped off. It reset okay.

At 8:22 PM, MIRF14 tripped off for the first of two times in fifteen minutes.

NuMI experts completed their target scans at 9:34 PM.

At 10:41 PM, BRF15 tripped off. It reset okay.

At 11:08 PM, BRF16 tripped off. It reset okay.

Monday April 4
The midnight shift began with Operations monitoring store 4074 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 4:58 AM, Operations reported that an Accumulator kicker had stopped working, but then five minutes later it resumed normal operations.

Operations terminated store 4074 at 5:03 AM.

Operations put the TeV into shot setup at 7:11 AM.

At 7:40 AM, the TeV suffered an abort. Operations held off resetting the abort so that experts could conduct an investigation.

The Plan for Monday
The plan for today is to establish a new store. This morning's combination store will be from the largest stack/stash ever recorded: the stack has 171.5E10 antiprotons, and the stash has 123.8E10 antiprotons.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/4/2005   email Fermilab